What is the answer of 5+y÷4=28÷4 an does it or does it not equal 23 an why?
Answer by idontknowdoyou542(2)

is your problem like this.? 5+ y/4 = 28/4??
... then if so,, multiply the problem to 4 to cancel out 4 in the denominator places..
4*(5+ y/4 = 28/4)
so the equation becomes like this..
20 + y = 28
then to find the value of y,, transpose 20 to the other side and change its sign.
y = 28-20
y= 8.. # no, it is not equal to 23..
# as to why,, I'm not sure but only a single value could satisfy the equation in which it is 7,, i forgot the name of those types of equation.
CHECk:
5 + 8/4 = 28/4
5+2= 28/4
4*(7)= 28
28=28.